OpenStudy (anonymous):

What is the vertex of the graph of y=-4(x+2)^2+5. Please help quickk!!!!!! I will give a medal for the right answer.

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png should be very conspicuous, the equation is already in "vertex form"

jigglypuff314 (jigglypuff314):

the vertex form of such a graph is y = a(x-h)^2 + k a = how much it is stretched/shrunk and (h , k) = vertex

OpenStudy (anonymous):

can u solve it ?

jigglypuff314 (jigglypuff314):

... there is nothing to be solved, just take numbers out of what was given to you y=-4(x+2)^2+5 ^ ^ y = a(x- h)^2+k when vertex is (h, k)

OpenStudy (anonymous):

here are the choices (2, 5) (-2, 5) (5, -2) (5, 2)

OpenStudy (anonymous):

Oh okay i understand.

jigglypuff314 (jigglypuff314):

but be careful that it is a Negative 2 comma 5 that you got... many people tend that make that mistake :)

OpenStudy (anonymous):

okay can you help with another Q

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

solve (x+2)^2=1

jigglypuff314 (jigglypuff314):

foil get all values to one side then factor

OpenStudy (anonymous):

can you show me how ?

jigglypuff314 (jigglypuff314):

okay, so do you know how to foil (x + 2)^2 = ?

OpenStudy (anonymous):

no

jigglypuff314 (jigglypuff314):

how about if it looked like this: (x + 2)(x + 2) = ?

jigglypuff314 (jigglypuff314):

(x + 2)^2 = (x + 2)(x + 2) = x^2 + 4x + 4 so you get x^2 + 4x + 4 = 1 - 1 -1 x^2 + 4x + 3 = 0 then factor so (x + 3)(x + 1) = 0 so solve for x when x + 3 = 0 and x + 1 = 0 that would give you your answers :)